# Emergent Coordination in Multi-Agent Language Models

### v1

@@ Acknowledgments
We thank Andrea Barolo for some early work on the implementation and Amritesh Anand for excellent research assistance on the project. We thank Fernando Rosas, Yonatan Belinkov, and Patrick Forber for comments. This work was supported by the D’Amore-McKim School of Business’ DASH initiative which provided access to crucial computational resources.
@@ A.1 Prompts
Persona Generation Generate list of 20 personas. each persona description should be about one paragraph long. personas should include diverse people with different backgrounds, jobs, skills, different preferences, and different personalities. format the output as follows ”You are [name], the [job title or professional identitiy]: [name] [rest of the persona]”
@@ A.4 Falsification Tests: Additional Method Details
We perform a falsification test of each of the three criteria introduced above in two different ways. In our robustness test, we remove time trends from the equal-share deviations via agent-level linear regression residualization to mitigate autocorrelation confounds. We then generate a null distribution of our statistics computed on column-wise block shuffled data (we use $B=200$ random shuffles for main analyses and confirm results using $B=1,000$ on the emergence criteria tests). We permute time indices jointly across agents in blocks of length $\ell$, preserving within-block identity alignment while disrupting cross-agent temporal coupling beyond short time frames. Finally, we compute the bias-corrected estimate as the observed value minus the median of the null. We then perform a Wilcoxon signed‑rank tests to test whether the bias‑corrected values exceed zero ($p(H1:>0)$. In some of our sensitivity tests we also perform a full within-row shuffle completely permuting guesses within each trial across agents (breaking agent identities), while preserving row constraints. This role differentiation or identity-specific coordination, including time-trend effects (tests identity-locked differentiation beyond task induced temporal dynamics). Comparison between results using the full row-shuffle vs. the column-block-shuffle separates signs of identity‑locked (specialized roles) from dynamic alignment (mutual adaptation, turn‑taking, oscillation). In addition to regression residualizing we also implement a functional baseline (see Appendix).
@@ A.5 Emergence Capacity: Full Result Details
Next, we test if there are signs of emergence capacity, again using both the p-value test on raw data and the test whether bias corrected values are above 0. We compute the dynamical synergy, which measures how much of the system’s future behavior is only predictable from the whole system and not any subset of its parts. The main test finds that about 32% of groups show significant emergence capacity with p-values below 0.05 (Plain: 37%; Persona: 44%; ToM 18%). A joint Fisher test of all p-values is also highly significant. As robustness test we check whether the bias corrected value of dynamical synergy is above 0 using the Wilcoxon test (Figure 2b). The test is highly significant overall, as well as in the Plain and Persona condition. The ToM condition is marginally significant ($p=0.099$).
@@ A.8 Robustness Tests using Three Bins
20% of individual experiment $p$-values are below 0.05 and Fisher test is highly significant (the largest $p$ value is in the ToM condition with $4.47\times 10^{-3}$ the others are much smaller). Wilcoxon signed rank test for $\mu>0$ of the bias-corrected and time-trend demeand data are as follows. Overall: $p=0.005$; Plain: $p=0.0004$; Persona: $p=3.2\times 10^{-6}$; ToM: $p=0.991$.
@@
A.9 Are Identities Persistent?
@@
A.10 Additional Details for Performance Analysis
What does synergy enable? In particular, does synergy affect performance? Connecting emergent properties of synergy and redundancy to performance is a challenging task for two main reasons. First, redundancy and synergy sit at opposite ends of the informational spectrum of correlations. Systems constrained by limited resources (such as energy or bandwidth) cannot maximize both simultaneously. The “informational budget” of correlations can be spent on shared overlap (redundancy) or on higher-order joint structures (synergy), but not both simultaneously (Williams & Beer, 2010; Varley et al., 2023; Luppi et al., 2024). Second, emergent properties are dichromatic—they require time to evolve (Bedau, 2002)—making them endogenous with run-length of the simulation. Namely, simulations in which the multi-agent system performs worse and takes longer to solve the task (or fails to solve the task altogether) provide them with more time to evolve sophisticated group coordination patterns; while successful runs terminate early, giving collectives less time to evolve emergent properties. This complicates meaningful analyses as it can make it appear that higher synergy is correlated with failure. In technical terms: time is an endogenous variable and confounds both performance and observed values of synergy and redundancy.
@@ A.10 Additional Details for Performance Analysis
A.11 Robustness Tests using “Early Synergy”
@@
A.12 Additional Results using Llama-3.1-8B
The majority of Llama agent groups fail to solve the task. In the vast majority of instances, groups get stuck in oscillating behavior of guessing too high (low) and then overcompensating by guessing too low (high), unable to break out of the cycle by establishing specialized differentiation across agents and pursuing complementary behavior.
Robustness tests using the emergence capacity measure are mixed. Using the emergence capacity test indicates that Llama agents can exhibit emergent synergy (10.5% of individual groups show $p$-values are below 0.05 compared to 32% of gpt-4.1; the joint Fisher test is highly significant). Robustness tests using bias corrected estimates on time-trend residualized data find significant signs of emergence in the Persona condition (Wilcoxon signed rank test $p=0.020$), marginal evidence in the ToM condition ($p=0.061$), and no evidence in the Plain condition ($p=0.252$; see Appendix for more).
@@
The test for emergence using the practical criterion shows no evidence for emergence (only 3.3% of individual groups show $p$-value below 0.05 and joint Fisher suggests no evidence with $p=1$ either overall or within any of the treatment conditions). The robustness test using the bias-corrected estimates on time-trend demeaned data suggest emergence in the Plain condition ($p=0.011$), no evidence in the Persona condition ($p=0.232$), and marginal evidence in the ToM condition ($p=0.068$).

### v2

@@ Acknowledgments
We thank Andrea Barolo for some early work on the implementation and Amritesh Anand for excellent research assistance on the project. We thank Fernando Rosas, Yonatan Belinkov, Patrick Forber, Koyena Pal, and David Bau for comments. This work was supported by the D’Amore-McKim School of Business’ DASH initiative which provided access to computational resources.
@@ A.1 Prompts
Persona Generation Generate a list of 20 personas. Each persona description should be about one paragraph long. Personas should include diverse people with different backgrounds, jobs, skills, different preferences, and different personalities. Format the output as follows ”You are [name], the [job title or professional identity]: [name] [rest of the persona].”
@@ A.4 Falsification Tests: Additional Method Details
We perform a falsification test of each of the three criteria introduced above in two different ways. In our robustness test, we remove time trends from the equal-share deviations via agent-level linear regression residualization to mitigate autocorrelation confounds. We then generate a null distribution of our statistics computed on column-wise block shuffled data (we use $B=200$ random shuffles for main analyses and confirm results using $B=1,000$ on the emergence criteria tests). We permute time indices jointly across agents in blocks of length $\ell$, preserving within-block identity alignment while disrupting cross-agent temporal coupling beyond short time frames. Finally, we compute the bias-corrected estimate as the observed value minus the median of the null. We then perform a Wilcoxon signed‑rank tests to test whether the bias‑corrected values exceed zero ($p(H1>0)$. In some of our sensitivity tests we also perform a full within-row shuffle completely permuting guesses within each trial across agents (breaking agent identities), while preserving row constraints. This role differentiation or identity-specific coordination, including time-trend effects (tests identity-locked differentiation beyond task induced temporal dynamics). Comparison between results using the full row-shuffle vs. the column-block-shuffle separates signs of identity‑locked (specialized roles) from dynamic alignment (mutual adaptation, turn‑taking, oscillation). In addition to regression residualizing we also implement a functional baseline (see Appendix).
@@ A.5 Emergence Capacity: Full Result Details
Next, we test if there are signs of emergence capacity, again using both the $p$-value test on raw data and the test whether bias corrected values are above 0. We compute the dynamical synergy, which measures how much of the system’s future behavior is only predictable from the whole system and not any subset of its parts. The main test finds that about 32% of groups show significant emergence capacity with $p$-values below 0.05 (Plain: 37%; Persona: 44%; ToM 18%). A joint Fisher test of all $p$-values is also highly significant. As robustness test we check whether the bias corrected value of dynamical synergy is above 0 using the Wilcoxon test (Figure 2b). The test is highly significant overall, as well as in the Plain and Persona condition. The ToM condition is marginally significant ($p=0.099$). See analyses using $K=3$ bins for additional discussion and evidence of significant synergy in the ToM condition.
@@ A.8 Robustness Tests using Three Bins
20% of individual experiment $p$-values are below 0.05 and Fisher test is highly significant (the largest $p$ value is in the ToM condition with $4.47\times 10^{-3}$ the others are much smaller). Wilcoxon signed rank test for $\mu>0$ of the bias-corrected and time-trend demeand data are as follows. Overall: $p=0.005$; Plain: $p=0.0004$; Persona: $p=3.2\times 10^{-6}$; ToM: $p=0.991$. Wilcoxon signed rank test for $\mu>0$ of the bias-corrected and functional-null demeand data are as follows. Overall: $p=0.948$; Plain: $p=0.999$; Persona: $p=0.896$; ToM: $p=0.025$. Using this stricter null indicates that only the ToM condition exhibits complementary cross-agent structure not explained by shared trends or independent feedback-following. That is, the stricter test isolates the construct of interest—synergy beyond coordination-free dynamics—and ToM is the only condition that passes it. The high $p$-value for ToM under the simpler time-trend demeaning should be taken as a methodological sensitivity (sparsity and a fat null). Together they show the ToM effect is robust to deconfounding but fragile under known low-power discretization. This is also in line with the stricter early synergy tests presented in Section A.12 which are also only passed by the ToM condition.
@@
A.9 Additional Results on Dynamical Mechanism
The main paper shows results for Total Stability: $I_{3}$ normalized by the entropy of the macro signal. Here we show results for $I_{3}$ as well (Figure A2). They are nearly identical given that the entropy of macro signal if $\approx 1$ given quantile binning with two bins.
Figure A2: $I_{3}$ mutual information (time-trend demean, bias corrected) with Wilcoxon >0>0 $p$-value indicators.
A.10 Are Identities Persistent?
@@
Together, the results suggest a nuanced picture of emergence in multi-agent systems. Persona and ToM produce agents with distinct, stable identities—moving agents on the differentiated-undifferentiated axis in Figure 1a. This differentiation shapes the space in which other agents can adapt, and specialization by one agent constrains the (useful) actions of the others (agents mutually constrain each other; cf. Tollefsen et al., 2013)—shifts on the interdependent-complementary axis. When combined with ToM, mutual adaptation both strengthens agent differentiation (creating the foundation for possible synergistic complementarity) while also increasing mutual alignment on shared goals (the target signal)—shifts on the integration axis. Differentiation enables synergy, but without redundancy and integration, synergy alone does not translate into higher collective performance. Agents in the ToM condition develop specialized, complementary roles while achieving stronger shared goal alignment: they are simultaneously more differentiated yet also more integrated along highly overlapping task-relevant information which the capacity for synergy can now exploit in productive ways.
A.11 Additional Details for Performance Analysis
What does synergy enable? In particular, does synergy affect performance? Connecting emergent properties of synergy and redundancy to performance is a challenging task for two main reasons. First, redundancy and synergy sit at opposite ends of the informational spectrum of correlations. Systems constrained by limited resources (such as energy or bandwidth) cannot maximize both simultaneously. The “informational budget” of correlations can be spent on shared overlap (redundancy) or on higher-order joint structures (synergy), but not both simultaneously (Williams and Beer, 2010; Varley et al., 2023; Luppi et al., 2024). Second, emergent properties are dichromatic—they require time to evolve (Bedau, 2002)—making them endogenous with run-length of the simulation. Namely, simulations in which the multi-agent system performs worse and takes longer to solve the task (or fails to solve the task altogether) provide them with more time to evolve sophisticated group coordination patterns; while successful runs terminate early, giving collectives less time to evolve emergent properties. This complicates meaningful analyses as it can make it appear that higher synergy is correlated with failure. In technical terms: time is an endogenous variable and confounds both performance and observed values of synergy and redundancy.
@@ A.11 Additional Details for Performance Analysis
A.12 Robustness Tests using “Early Synergy”
@@
A.13 Experiments with other Models
To test generalizability of our findings, we conducted experiments using four other models:
  1. 1.
    Llama 3.1 8B,
  2. 2.
    Llama 3.1 70B Instruct,
  3. 3.
    Google Gemini 2.0 flash, and
  4. 4.
    Qwen3 235B A22B Instruct 2507.
We conducted 100 replications of the group experiment with ten agents and temperature of 1.0 of each model and each treatment condition.
Before turning to comparative results, we first address a structural challenge that emerged with Qwen3—one that speaks to a broader issue with deploying reasoning models in multi-agent coordination settings.
As a reasoning model, Qwen3 exhibited persistent looping behavior during its chain-of-thought process—a known failure mode of reasoning models (Pipis et al., 2025). These models are particularly susceptible to self-reflection and deductive reasoning traps (Baek and Tegmark, 2025). In our setting, we identify three specific patterns of oscillatory reasoning that arise reliably as soon as group feedback becomes inconsistent with the individual binary search strategy.
Adjacent Integers.
When group feedback places an agent between two adjacent integers, the agent fails to recognize that the ambiguity of that feedback is driven by the unobservability of other agents’ guesses. A typical reasoning pattern was: “8 was too low but 9 was too high, but guesses need to be integers, let me re-read the instructions …”). The agent interprets the goal literally and realizes they are mutually incompatible, not realizing that changes in the guesses of other agents explain the seeming inconsistent feedback.
Dual Strategies.
Agents fail to reconcile the strategy inconsistency of both reasoning about the task (applying binary search strategy) while also coordinating with other agents. A typical reasoning pattern was: “I should guess the midpoint 5 … but I also need to adjust to others … so maybe I should guess higher … but that’s not the midpoint … let me re-read the instructions …”.
Mutual Mental Modeling.
We observe four different categories of mutual mental modeling traps, each representing a distinct cognitive stance.
  1. 1.
    Uncertainty Acknowledgment: “Perhaps the other players guess is changing. But no information.” Agents recognize that other agents may be changing their behavior and admit epistemic helplessness. This shows awareness of others, without actual modeling.
  2. 2.
    Causal Attribution Under Ambiguity: “Perhaps the mystery number is for the sum, and when you guess 10, the sum was too high, but maybe the other guesses changed.” Here, the agent is using mental modeling to interpret the group feedback. This shows mutual modeling as a diagnostic tool to explain the inconsistency.
  3. 3.
    Strategic Simplification: “Given the instruction to use binary search, I think we must assume others are constant.” This is a meta-level decision to collapse the modeling problem by deliberately treating other agents as static to make the task tractable. In some sense, this is the most sophisticated move as it recognizes that mutual modeling creates infinite regress and chooses a pragmatic simplification.
  4. 4.
    Concrete Joint-State Estimation: “I should guess 9, and others 8 — sum 17. So guess 9. Or if others are at 8, I should guess 9. But I don’t know.” This is an actual attempt to simulate the joint outcome by assigning specific values to other agents’ guesses yet it dissolves into uncertainty.
These examples illustrate that the model is aware of the task ambiguity and often correctly recognizing the oscillating pattern (often, reasoning trace includes the word phase “[we are] oscillating”). That is, the failure is not due to reasoning error. Instead it is failure to terminate under unresolved ambiguity. We term this paralysis under coordination ambiguity. Notably, the paralysis persists even at a high temperature setting ($T=1.0$), which should increase stochastic exploration and disrupt deterministic decoding loops, indicating that the instability arises from internal reasoning dynamics rather than sampling artifacts. As a practical remedy, we managed to improve robustness by adding one extra line to the prompt to to the Qwen experiments:
[…]
If you find yourself re-evaluating your guess more than once, repeat your last guess.
[…]
@@ A.13 Experiments with other Models
Overall, we find evidence of multi-agent LLM system capacity for emergence in other models, with some variation across models with different capacity and treatment condition (Table A2). We find the task is quite hard for Llama 3.1 8B, and the majority of groups fail to solve it. In the vast majority of instances, groups get stuck in oscillating behavior of guessing too high (low) and then overcompensating by guessing too low (high), unable to break out of the cycle by establishing specialized differentiation across agents and pursuing complementary behavior. We find evidence of emergence using emergence capacity test (10.5% of individual groups show $p$-values are below 0.05; the joint Fisher test is highly significant). Robustness tests using bias corrected estimates on time-trend residualized data find significant signs of emergence in the Persona condition but marginal and no evidence in the two other conditions. The test for emergence using the practical criterion shows no evidence for emergence (only 3.3% of individual groups show $p$-value below 0.05 and joint Fisher suggests no evidence with $p=1$ but the robustness test using the bias-corrected estimates on time-trend demeaned data suggest emergence in the Plain condition ($p=0.011$), no evidence in the Persona condition ($p=0.232$), and marginal evidence in the ToM condition ($p=0.068$).
Groups of Llama 3.1 70B agents perform substantially better. We find strong evidence of emergence using the emergence capacity criterion (both overall and in the persona condition). All tests with the practical criterion are significant. Groups of Gemini 2.0 flash agents perform even higher. We find strong evidence of emergence using the emergence capacity criterion (overall Fisher test is significant, as well es significant robustness test in the Plain and ToM condition). Performance of groups of Qwen3 agents falls between Llama 70B and Gemini 2.0 flash. We find strong evidence of emergence using both the emergence capacity criterion and the practical criterion with most tests showing highly significant $p$-values (only the Wilcoxon test in the Plain is marginal at 0.052). The insignificant $I_{3}$ in groups of Qwen agents could be a result of the failure modes outlined above: the reasoning process of alignment with other agents is never successful given the paralysis under coordination ambiguity. The insignificant $G_{3}$ in groups of Llama 3.1 8B agents could be a result of that model’s generally much lower ToM ability.
Llama-3.1-8B Llama-3.1-70B Gemini 2.0 flash Qwen3
Group Performance
SuccessPlain 11% 53% 60% 51%
SuccessPersona 14% 59% 75% 71%
SuccessToM 5.5% 61% 71% 58%
Emergence Capacity
Fisher 0.000 0.000 0.000 0.000
[…]
Practical Criteria
[…]
WilcoxonPlain 0.011 0.002 0.215 0.052
[…]
Groups with Significant Differentiation
Plain 86% 68% 64% 89%
Persona 92% 77% 77% 94%
ToM 80% 86% 86% 97%
I3
Fisher 0.000 0.000 0.000 1.000
G3
Fisher 1.000 0.000 0.000 0.000
Table A2: Experiment results with other LLM models.
A.14 Specific Persona Effects
Both the Persona and ToM conditions rely on personas. Do the specific personas—or combinations of personas—present in a multi-agent system affect performance or the emergence of higher-order structures? We explore these questions with several test. Recall that our experiments are based on 20 different personas that are generated once globally, and then reused throughout the experiments by sampling 10 persona from the library (without replacement). First, we perform joint F-tests (based on linear regression) of the null hypothesis that all persona dummies have no effect on the outcome (i.e., no specific persona substantially affects group outcomes and all coefficients equal zero). Rejecting the null would indicate that at least one of the persona indicators is significantly associated with the outcome variable. We find no evidence that the presence of specific personas effect performance ($F=1.217$; $p=0.239$), number of rounds to success ($F=1.356$; $p=0.144$), PID synergy ($F=0.839$; $p=0.66$), or PID redundancy ($F=0.596$; $p=0.910$). These results suggest that what matters is the presence of personas overall—markers that allow the LLM agents to center their differentiation on—rather than who those personas are specifically.
Next, we used an LMRA (Eloundou et al., 2024; Riedl and Weidmann, 2025) to judge the pairwise similarity of the personas in terms of their personality and background (we used gpt-oss-120b to assess similarity on a 0-10 scale). Even though judgments are subjective, the input they judge is standardized, controlled, information-rich, and deliberately designed to reveal deep-level characteristics in line with personas used in vignette studies (Aguinis and Bradley, 2014). This allows the LMRA to integrate many deep-level characteristics in a controlled and comparable format. From these, we compute group-level diversity as one minus the mean pairwise similarity rating, following the dispersion-based operationalization of separation diversity in Harrison and Klein (2007) (mean diversity: 5.457; SD = 0.385). In the last step, we explore whether group diversity is systematically related with key group-level outcomes using linear regression. We find no effect of diversity on performance ($\beta=0.067$; $p=0.403$), number of rounds to success ($\beta=-0.195$; $p=0.877$), PID synergy ($\beta=-0.000$; $p=0.856$), or PID redundancy ($\beta=0.003$; $p=0.113$). These results again suggest that there are no strong effects for specific combinations of personas.